# Unraveling Indirect In-Context Learning Using Influence Functions

### v2

@@ 4.1 Experimental Setup
Evaluation Data. For Mixture of Tasks, we collect a generalized pool of examples from different tasks such that the input $x$ and output $y$ pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ do not necessarily correspond to the same task as the test input $x_{\text{test}}$. The evaluation task pool contains three samples each from 28 different tasks from MMLU Hendrycks et al. (2020), BigBench Srivastava et al. (2022), StrategyQA Geva et al. (2021) and CommonsenseQA Talmor et al. (2018). We evaluate the ICL accuracy, using this train set, on 12 different tasks from MMLU and BigBench.
For Noisy ICL, we employ the noisy dataset framework from Kwon et al. (2023). In their work, the four binary classification GLUE datasets Wang (2018) MRPC, QQP, QNLI, and SST2 are utilized. To simulate a scenario where a portion of the data samples are noisy, 20% of the training data samples are randomly selected and their labels are flipped. We use these noisy datasets as the candidate pool in our experiments and evaluate the ICL accuracy.
Base LLM. In Mixture of Tasks, for $k=3$ shots, we conduct ICL experiments on Llama-2-13b-chat Touvron et al. (2023), Mistral-7b-v0.3 Jiang et al. (2023), Zephyr-7b-beta Tunstall et al. (2023) and Qwen2.5-3b Team (2024). For $k=5$ shots we conduct experiments on Llama-2-13b-chat. We extend on the framework designed by Gupta et al. (2023a, b). The temperature is set to 0 for inference. For Noisy ICL, we conduct experiments on Llama-2-13b-chat for mislabeled data detection and Llama-3-8b for backdoor defense. All of our experiments run on 8$\times$×NVIDIA RTX 6000 Ada GPUs.
@@ 4.2 Method and Baseline Configurations
Mixture of Tasks. We construct 4 IF-only methods. 2 based on the Surrogate Model based approach, Sur and 2 based on the Pretrained LLM weights based approach, Pre. We test Data-Inf and TracIn based versions of these approaches, namely, Surrogate Model-DataInf SurD, Surrogate Model-TracIn SurT, Pretrained Model-DataInf PreD and Pretrained Model-TracIn PreT. As mentioned before, SurD and SurT use RoBERTa as the surrogate model, whereas PreD and PreT use Llama2-13b-chat as the pretrained LLM.
@@
[…]
For baselines, we again compare our IF based approaches with BSR, Cosine Similarity and BM25.
@@
4.3 Results on Mixture of Tasks
We present the results on Mixture of Tasks in Figure 2. Additional results for varying the number of shots ($k$) and multiple LLMs are provided in Appendix C.1. Further, results for the alternative TracIn IF method are provided in Appendix C.2. Results for Pretrained Gradients combined with BertScore and Cosine Similarity are presented in Appendix C.3. We also present results on using DeBERTa He et al. (2021a, b) as an alternative surrogate model in Appendix C.4. Finally, we present results on Qwen2.5-3b Team (2024) in Appendix C.5.
@@ 4.3 Results on Mixture of Tasks
Surrogate Models outperform Pretrained Gradients. We see that surrogate models outperformed Pretrained Gradients in demonstration selection for the Mixture of Tasks setting in both the $k=3$ and $k=5$ shots. The fine-tuning of the surrogate model leads it to better capture the test task affinity of the demonstration pool.
DataInf is better than TracIn as an IF method. The speed gains of TracIn come at a cost of performance as the DataInf method of IF computation routinely outperformed TracIn. TracIn likely underperforms because it does not utilize critical second order gradient information since the Hessian $H(\theta^{*})$ is assumed to be the identity matrix. This trend has also been observed in past work on IF methods Chhabra et al. (2024a).
@@ 4.4 Results on Noisy ICL
Pruning hurts not helps performance. We can see that pruning actively hurts performance as Figure 3 shows that all 3 types of BertScore pruning and all 3 types of Cosine pruning had lower average scores than BertScore and Cosine Similarity. This might be due to the fact that we are removing potentially helpful samples from the demonstration pool, even if they might have noisy labels.
Figure 3: Average performance of the baselines across the 4 datasets.
We further provide results on varying the noise levels in the datasets in D.1, varying the hyperparameters we tested in D.2 and an experiment analyzing the effectiveness of IF’s in detecting Noisy ICL in D.3.
@@ 4.5 Computational Complexity
We present the worst case time complexity (for inference) for our methods and related baselines in Table 2. As can be observed, our methods are comparable, if not more efficient than the other baselines. Note that the Sur methods require an additional fine-tuning step on a smaller surrogate model before the gradients are extracted, which the Pre methods do not. Furthermore, note that TracIn as an influence method is much faster than Hessian-based approaches (e.g. DataInf) as it assumes that the Hessian is the identity matrix. While this leads to more efficient influence computation, it comes at the cost of lower estimation performance, as our results with TracIn also show. Additionally, we present the maximum GPU memory consumption while performing demonstration selection in Appendix E, scalability to large models in Appendix F and scalability to large datasets in Appendix G
Table 2: Computational complexity for each test sample at inference where $N$ is #demonstration samples, $p$ is #model parameters, $d$ is embedding size, $K$ is the max length among all candidates, $L$ is the length (in tokens) of the test input, $Z$ is #ngrams
@@ 5 Related Work
In-Context Learning (ICL). Following the scaling of model sizes and learning resources (Mann et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023), LLMs have gained emergent abilities for efficient inference-time adaptation via ICL Mann et al. (2020). However, ICL is critically sensitive to demonstration pool examples (An et al., 2023; Liu et al., 2021a; Zhang et al., 2022) and selection strategies (Rubin et al., 2021; Mavromatis et al., 2023). One line of work studies example scoring and retrieval, utilizing model-agnostic heuristic metrics like perplexity (Gonen et al., 2022), mutual information (Sorensen et al., 2022), semantic similarity (Liu et al., 2021a; Gupta et al., 2023b), etc. to select demonstrations. Another line of work optimizes selection based on empirically verified desirable features a priori, e.g. diversity (Su et al., 2022; Ye et al., 2023), coverage (Gupta et al., 2023a), etc. However, prior work assumes that the demonstration distribution is aligned with task distribution, which is not always the case (Chatterjee et al., 2024). Our work serves as a first to investigate ICL demonstration selection in the task and dataset quality shifts in the ICL settings.
@@
Influence Functions. Influence functions (IFs) comprise a set of methods from robust statistics Hampel (1974); Cook and Weisberg (1982) that have been recently proposed for deep learning data valuation and can provide a conceptual link that traces model performance to samples in the training set. For gradient-based models trained using empirical risk minimization, IFs can be used to approximate sample influence without requiring actual leave-one-out retraining. For deep learning models, the seminal work by Koh and Liang (2017) utilized a Taylor-series approximation and LiSSA optimization (Agarwal et al., 2017) to compute sample influences. Follow-up works such as Representer Point (Yeh et al., 2018) and Hydra (Chen et al., 2021) sought to improve IF performance for deep learning models, constrained to vision applications. More recently, efficient influence estimation methods such as DataInf (Kwon et al., 2023), Arnoldi iteration (Schioppa et al., 2022), and Kronecker-factored approximation curvature (Grosse et al., 2023) have been proposed which can be employed for larger generative language models, such as LLMs. Some other simpler IF approaches simply consider the gradients directly as a measure of influence (Pruthi et al., 2020; Charpiat et al., 2019), followed by some ensemble strategies (Bae et al., 2024; Kim et al., 2024). Recent work has also found that self-influence only on the training set can be a useful measure for detecting sample influence (Bejan et al., 2023; Thakkar et al., 2023).
@@
IFs have been utilized with great success in a number of application scenarios (e.g. classification Chhabra et al. (2024a); Koh and Liang (2017), generative models Kwon et al. (2023); Schioppa et al. (2022); Grosse et al. (2023), active learning Chhabra et al. (2024b); Liu et al. (2021b), etc.). Moreover, while some recent works have considered using influence for selecting direct demonstrations (Nguyen and Wong, 2023; Van et al., 2024), neither of them has considered their effect on inductive bias selection in the indirect ICL setting, which is the focus of our work.
@@ 6 Conclusion
Limitations
The main limitation of Influence Functions is that they are costly to compute, especially for large datasets and LLMs with lots of parameters. This is why we opted to fine-tune a smaller model such as RoBERTa and use pretrained LLMs for our methods. Further performance gains can be attained at the cost of computational speed if fine-tuned LLMs are employed instead. As research on influence estimation methods for LLMs is currently ongoing, faster influence functions developed in the future can also be utilized with our methods for highly efficient and accurate ICL performance.
@@ C.4 Results using a different surrogate model
Here we present results on Qwen2.5-3b Team (2024) for $k=3$ shots on the Mixture of Tasks setting. We compare BSR with our two IF based BSR methods Sur[D,BSR] and Pre[D,BSR]. We see in Table 16 that for an advanced model like Qwen2.5-3b, the unfinetuned pretrained gradients slightly outperform the finetuned gradients from the RoBERTa model.
@@
C.6 Ablation analysis on the benefits of combining IF with semantic similarity methods
@@ Appendix F Scalability to Large Models
We compare the time it takes to extract test set gradients and compute influence scores. We compute IF via the DataInf method, comparing RoBERTa-Large (125 million parameters), Llama2-13b-chat (13 billion parameters), and Llama2-70b-chat (70 billion parameters) on the MMLU-moral-disputes dataset with 200 test samples.
@@ Appendix G Scalability to large datasets
[…]
@@ Appendix H IF-based demonstrations as a backdoor defense strategy to mitigate backdoor attacks
We use our Sur[D,BSR] strategy to the SST-2 dataset Wang (2018), which has been poisoned with two distinct forms of backdoor attacks: (1) AddSent Dai et al. (2019), where a semantic trigger is introduced by inserting an innocuous sentence (e.g., “I watched this 3D movie last weekend”), and (2) Style Qi et al. (2021), a style-based attack that performs backdoor poisoning through text style transfer (e.g., transforming text into a Biblical style).
We evaluate the Attack Success Rate (ASR), defined as the percentage of non-target label test instances that are misclassified as the target label when evaluated on a poisoned dataset. For comparison, we consider a task-aware demonstration selection baseline Diao et al. (2023), which selects 5 demonstrations from a clean, task-specific dataset. All experiments are conducted using the LLaMA3-8B model Grattafiori et al. (2024). We present our results in Table 23.

### v3

@@ 4.1 Experimental Setup
Evaluation Data. For Mixture of Tasks, we collect a generalized pool of examples from different tasks such that the input $x$ and output $y$ pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ do not necessarily correspond to the same task as the test input $x_{\text{test}}$. The evaluation task pool contains three samples each from 28 different tasks from MMLU (Hendrycks et al., 2020), BigBench (Srivastava et al., 2022), StrategyQA (Geva et al., 2021) and CommonsenseQA (Talmor et al., 2018). We evaluate the ICL accuracy, using this train set, on 12 different tasks from MMLU and BigBench.
For Noisy ICL, we employ the noisy dataset framework from Kwon et al. (2023). In their work, the four binary classification GLUE datasets (Wang, 2018) MRPC, QQP, QNLI, and SST2 are utilized. To simulate a scenario where a portion of the data samples are noisy, 20% of the training data samples are randomly selected and their labels are flipped. We use these noisy datasets as the candidate pool in our experiments and evaluate the ICL accuracy.
[…]
@@ 4.2 Method and Baseline Configurations
Mixture of Tasks. We construct 4 IF-only methods. 2 based on the Surrogate Model based approach, Sur and 2 based on the Pretrained LLM weights based approach, Pre. We test Data-Inf and TracIn based versions of these approaches, namely, Surrogate Model-DataInf SurD, Surrogate Model-TracIn SurT, Pretrained Model-DataInf PreD and Pretrained Model-TracIn PreT. As mentioned before, SurD and SurT use RoBERTa as the surrogate model, whereas PreD and PreT use Llama2-13b-chat as the pretrained LLM. Additionally, we test traditional semantic approaches, such as BSR and Cosine Similarity (Cos), as well as retrieval based approaches, such as BM25, as baselines. Finally, we test the combination of the aforementioned traditional and IF methods as well.Specifically, we denote these as:
$\bm{\textbf{{Model}}}_{[\bm{\textbf{{If}}},\bm{\textbf{{Sel}}}]},\text{ where }\bm{\textbf{{Model}}}\in\{\bm{\textbf{{Sur}}},\bm{\textbf{{Pre}}}\},\bm{\textbf{{If}}}\in\{\bm{\textbf{D}},\bm{\textbf{T}}\},\text{ and }\bm{\textbf{{Sel}}}\in\{\bm{\textbf{{Cos}}},\bm{\textbf{{BSR}}}\}.$
@@
4.3 Results on Mixture of Tasks
@@
We present the results on Mixture of Tasks in Figure 2. Additional results for varying the number of shots ($k$) and multiple LLMs are provided in Appendix C.1. Further, results for the alternative TracIn IF method are provided in Appendix C.2. Results for Pretrained Gradients combined with BertScore and Cosine Similarity are presented in Appendix C.3. We also present results on using DeBERTa (He et al., 2021a; b) as an alternative surrogate model in Appendix C.4. Finally, we present results on Qwen2.5-3b (Team, 2024) in Appendix C.5 and results on Llama-3-70b (Grattafiori et al., 2024) in Appendix C.6.
Figure 3: Average performance of the baselines across the 4 datasets.
@@
Surrogate Models outperform Pretrained Gradients. We see that surrogate models outperformed Pretrained Gradients in demonstration selection for the Mixture of Tasks setting in both the $k=3$ and $k=5$ shots. The fine-tuning of the surrogate model leads it to better capture the test task affinity of the demonstration pool. However, it is interesting to note that on more recent LLMs Appendix C.5 and LLMs with larger number of parameters Appendix C.6 PreTrained Gradients outperformed the surrogate models.
DataInf is better than TracIn as an IF method. The speed gains of TracIn come at a cost of performance as the DataInf method of IF computation routinely outperformed TracIn. TracIn likely underperforms because it does not utilize critical second order gradient information since the Hessian $H(\theta^{*})$ is assumed to be the identity matrix. This trend has also been observed in past work on IF methods (Chhabra et al., 2025).
@@ 4.4 Results on Noisy ICL
Pruning hurts not helps performance. We can see that pruning actively hurts performance as Figure 3 shows that all 3 types of BertScore pruning and all 3 types of Cosine pruning had lower average scores than BertScore and Cosine Similarity. This might be due to the fact that we are removing potentially helpful samples from the demonstration pool, even if they might have noisy labels. We further provide results on varying the noise levels in the datasets in D.1, varying the hyperparameters we tested in D.2 and an experiment analyzing the effectiveness of IF’s in detecting Noisy ICL in D.3.
Table 2: Computational complexity for each test sample at inference. NN is #demonstration samples, $p$ is #model parameters, $d$ is embedding size, $K$ is the max candidate length, $L$ is the length (in tokens) of the test input, $Z$ is #ngrams
IF protect against adversarial noise. The results presented in Appendix H show that the IF-based indirect-ICL paradigm mitigates backdoor attacks, reducing ASR by 32.89% on average. Even without task-specific data, demonstrations guided by a model’s inductive bias offer a strong task-agnostic backdoor defense.
@@ 4.5 Computational Complexity
We present the worst case time complexity (for inference) for our methods and related baselines in Table 2. As can be observed, our methods are comparable, if not more efficient than the other baselines. Note that the Sur methods require an additional fine-tuning step on a smaller surrogate model before the gradients are extracted, which the Pre methods do not. Furthermore, note that TracIn as an influence method is much faster than Hessian-based approaches (e.g. DataInf) as it assumes that the Hessian is the identity matrix. While this leads to more efficient influence computation, it comes at the cost of lower estimation performance, as our results with TracIn also show. Additionally, we present the maximum GPU memory consumption while performing demonstration selection in Appendix E, scalability to large models in Appendix F and scalability to large datasets in Appendix G. The experiments demonstrate the feasibility of applying our methods at increasing scales. Nevertheless, it is important to recognize that improving the efficiency of IFs remains an active research area, and such advancements can be readily integrated into our approach to further enhance the efficiency and accuracy of Indirect-ICL performance.
@@ 5 Related Work
In-Context Learning (ICL). Following the scaling of model sizes and learning resources (Mann et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023), LLMs have gained emergent abilities for efficient inference-time adaptation via ICL (Mann et al., 2020). However, ICL is critically sensitive to demonstration pool examples (An et al., 2023; Liu et al., 2021a; Zhang et al., 2022) and selection strategies (Rubin et al., 2021; Mavromatis et al., 2023). One line of work studies example scoring and retrieval, utilizing model-agnostic heuristic metrics like perplexity (Gonen et al., 2022), mutual information (Sorensen et al., 2022), semantic similarity (Liu et al., 2021a; Gupta et al., 2023b), etc. to select demonstrations. Another line of work optimizes selection based on empirically verified desirable features a priori, e.g. diversity (Su et al., 2022; Ye et al., 2023), coverage (Gupta et al., 2023a), etc. However, prior work assumes that the demonstration distribution is aligned with task distribution, which is not always the case (Chatterjee et al., 2024). Our work serves as a first to investigate ICL demonstration selection in the task and dataset quality shifts in the ICL settings.
@@
Influence Functions. Influence functions (IFs) comprise a set of methods from robust statistics (Hampel, 1974; Cook & Weisberg, 1982) that have been recently proposed for deep learning data valuation and can provide a conceptual link that traces model performance to samples in the training set. For gradient-based models trained using empirical risk minimization, IFs can be used to approximate sample influence without requiring actual leave-one-out retraining. For deep learning models, the seminal work by Koh & Liang (2017) utilized a Taylor-series approximation and LiSSA optimization (Agarwal et al., 2017) to compute sample influences. Follow-up works such as Representer Point (Yeh et al., 2018) and Hydra (Chen et al., 2021) sought to improve IF performance for deep learning models, constrained to vision applications. More recently, efficient influence estimation methods such as DataInf (Kwon et al., 2023), Arnoldi iteration (Schioppa et al., 2022), and Kronecker-factored approximation curvature (Grosse et al., 2023) have been proposed which can be employed for larger generative language models, such as LLMs. Some other simpler IF approaches simply consider the gradients directly as a measure of influence (Pruthi et al., 2020; Charpiat et al., 2019), followed by some ensemble strategies (Bae et al., 2024; Kim et al., 2024). Recent work has also found that self-influence only on the training set can be a useful measure for detecting sample influence (Bejan et al., 2023; Thakkar et al., 2023).
@@
IFs have been utilized with great success in a number of application scenarios (e.g. classification (Chhabra et al., 2025; Koh & Liang, 2017), generative models (Kwon et al., 2023; Schioppa et al., 2022; Grosse et al., 2023), active learning (Chhabra et al., 2024; Liu et al., 2021b), layer-quality estimation (Askari et al., 2025), etc.). Moreover, while some recent works have considered using influence for selecting direct demonstrations (Nguyen & Wong, 2023; Van et al., 2024), neither of them has considered their effect on inductive bias selection in the indirect ICL setting, which is the focus of our work.
@@ 6 Conclusion
Reproducibility Statement
[…]
@@ C.4 Results using a different surrogate model
Here we present results on Qwen2.5-3b (Team, 2024) for $k=3$ shots on the Mixture of Tasks setting. We compare BSR with our two IF based BSR methods Sur[D,BSR] and Pre[D,BSR]. We see in Table 16 that for an advanced model like Qwen2.5-3b, the unfinetuned pretrained gradients slightly outperform the finetuned gradients from the RoBERTa model.
@@
C.6 Results on Llama-3-70b
Here’s an explanation for the Llama-3-70B results following the style of your previous explanation:
Here we present results on Llama-3-70B for $k=3$ shots on the Mixture of Tasks setting. We compare both BSR and Cosine similarity baselines with our IF-based methods: Sur[D,BSR], Pre[D,BSR], Sur[D,Cos], and Pre[D,Cos]. We see in Table 17 that for a large-scale model like Llama-3-70B, the pretrained gradients combined with cosine similarity (Pre[D,Cos]) achieve the highest average performance at 83.256%. Notably, the pretrained-gradient based methods outperform the surrogate-model based methods in a pattern similar to the Qwen2.5-3b results. This suggests that the pretrained gradients in the newer-larger models are more representative of model’s inductive bias than the older-smaller models.
Table 17: Performance comparison between BSR, Sur[D,BSR], Pre[D,BSR], Cos, Sur[D,Cos], and Pre[D,Cos] for Llama-3-70B.
Dataset BSR Sur[D,BSR] Pre[D,BSR] Cos Sur[D,Cos] Pre[D,Cos] medical-genetics 95 94 97 97 98 98 prof-psychology 92.5 91 90.5 91.5 91.5 90.5 formal-logic 72.22 77.78 74.6 73.81 73.02 79.37 moral-disputes 89 91 90 87 90.5 91 public-relations 86.36 84.55 84.55 83.64 85.45 86.36 comp-security 86 86 86 87 88 87 astronomy 98.68 98.68 97.37 98.03 97.37 99.34 abstract-algebra 68 66 67 65 65 66 nutrition 95.5 93.5 93.5 95.5 95.5 94 high-school-biology 93.5 94 93.5 94 94.5 94 formal-fallacies 59 66.5 60.5 59.5 63 66 tracking-3 43.5 46 47 48 45.5 47.5 Average 81.605 82.418 81.793 81.665 82.278 83.256
C.7 Ablation analysis on the benefits of combining IF with semantic similarity methods
@@ Appendix F Scalability to Large Models
We compare the time it takes to extract test set gradients and compute influence scores. We compute IF via the DataInf method, comparing RoBERTa-Large (125 million parameters), Llama2-13b-chat (13 billion parameters), and Llama2-70b-chat (70 billion parameters) on the MMLU-moral-disputes dataset with 200 test samples. The results are provided in Table 22.
@@ Appendix G Scalability to large datasets
Finally, in MoT, the computational time of IF can further be optimized by only computing IF for the 2kk shots being pruned by BSR or Cosine similarity instead of the entire set of training demonstrations. Another optimization to the DataInf code could be replacing their handling of gradients with tensor operations instead of the current dict of dicts format. This enables the use of GPU processing for influence computation instead of CPU and can offer a considerable runtime speedup.
@@ Appendix H IF-based demonstrations as a backdoor defense strategy to mitigate backdoor attacks
We use our Sur[D,BSR] strategy to the SST-2 dataset (Wang, 2018), which has been poisoned with two distinct forms of backdoor attacks: (1) AddSent (Dai et al., 2019), where a semantic trigger is introduced by inserting an innocuous sentence (e.g., “I watched this 3D movie last weekend”), and (2) Style (Qi et al., 2021), a style-based attack that performs backdoor poisoning through text style transfer (e.g., transforming text into a Biblical style).
We evaluate the Attack Success Rate (ASR), defined as the percentage of non-target label test instances that are misclassified as the target label when evaluated on a poisoned dataset. For comparison, we consider a task-aware demonstration selection baseline (Diao et al., 2023), which selects 5 demonstrations from a clean, task-specific dataset. All experiments are conducted using the LLaMA3-8B model (Grattafiori et al., 2024). We present our results in Table 24.